# ExoIris: fast exoplanet transmission spectroscopy in Python

### v1

@@ 4 Discussion
A key strength of ExoIris lies in its explicit parameterization of wavelength-dependent quantities, such as the planet-star radius ratio and limb darkening coefficients, using spline models. This approach decouples the spectral resolution of the transmission spectrum from the resolution of the data, allowing users to balance model complexity against scientific goals. In addition, the inclusion of group-based structures for instrumental offsets, noise properties, and transit timing variations enables accurate joint modeling of heterogeneous datasets.
The example analysis of WASP-39 b highlights the effectiveness of these methods. ExoIris accommodates diverse instrument modes, detector configurations, and data characteristics within a single modeling framework. The results reproduce the transmission spectrum derived by Carter et al. (2024) while also providing improved flexibility for sensitivity testing, including the impact of systematic offsets and timing discrepancies.
I acknowledge support by the Spanish Ministry of Science and Innovation with the Ramon y Cajal fellowship number RYC2021-031798-I. Funding from the University of La Laguna and the Spanish Ministry of Universities is acknowledged.

### v2

@@ IV Discussion
A key strength of ExoIris lies in its explicit parameterization of wavelength-dependent quantities, such as the planet-star radius ratio and limb darkening, using interpolating models. This approach decouples the spectral resolution of the transmission spectrum from the data resolution, allowing users to balance model complexity against scientific goals. In addition, the inclusion of group-based structures for instrumental offsets, noise properties, and transit timing variations enables accurate joint modeling of heterogeneous datasets.
The example analysis of WASP-39 b highlights the effectiveness of these methods. ExoIris accommodates diverse instrument modes, detector configurations, and data characteristics within a single modeling framework. The results reproduce the transmission spectrum derived by Carter et al. (2024) and provide improved flexibility for sensitivity testing, including the impact of systematic offsets and timing discrepancies.
Additionally, ExoIris is computationally efficient. The 24 WASP-39 b joint analyses ($n_{\mathrm{k}}$ of 25, 60, 130, and 230, repeated for nearest-neighbor, linear, quadratic B-spline, cubic B-spline, PCHIP, and Makima radius-ratio interpolation) were all carried out during a single day using a MacBook Pro with an M3 Max CPU. High computational efficiency combined with the ability to perform iterative and branching workflows allows for an exploratory approach where the users can test hypotheses and refine models dynamically.
Finally, ExoIris remains under active development, driven by community feedback. Future updates will focus on further increasing the adaptability of the interpolation-based transmission spectrum model and improving the efficiency of the optimization and posterior sampling routines.
I acknowledge support by the Spanish Ministry of Science and Innovation with the Ramón y Cajal fellowship number RYC2021-031798-I. Funding from the University of La Laguna and the Spanish Ministry of Universities is acknowledged.